If t1andt2 are the times of flight for two complimentary angles then the range of projectile R is.???

Grade:11

3 Answers

Ravneet
5 years ago
complimentary angles :   two angles whose sum is 90°.
RANGE OF PROJECTILE   =      ( u2 sin2x)/g            where x = angle 
so sin2x = 2sinxcosx if angles are complimentry ,say , a and 90-a 
sin2a = 2sinacosa 
sin2(90-a) = 2sin(90-a)cos(90-a) = 2cosasina
Now t1 and t2 be two time of flight 
As sin2x is same , u and g are constant 
so t1 = t2  
 
Arun
5 years ago
Dear student
 
R = u² sin2A /g
 
Now
 
R = u² sin 2(90-A) /g
 
R = u² sin 2A/g
 
Same Range
 
Hence of a particle is projected woth two complimentary angles, its range is same
Ravneet
5 years ago
it means the range is same for complementary angles 
range for t1 = range for t2.
example : 

For complementary angles, the two possible sets could be -

θ1 = 30 deg , θ2 = 60 deg; or

θ1 = 60 deg , θ2 = 30 deg

For all of the above angles of projection, sin 2θ comes out to be the same value